### v2

# Attribute or Abstain: Large Language Models as Long Document Assistants

Abstract
LLMs can help humans working with long documents, but are known to hallucinate. Attribution can increase trust in LLM responses: The LLM provides evidence that supports its response, which enhances verifiability. Existing approaches to attribution have only been evaluated in RAG settings, where the initial retrieval confounds LLM performance. This is crucially different from the long document setting, where retrieval is not needed, but could help. Thus, a long document specific evaluation of attribution is missing. To fill this gap, we present LAB, a benchmark of 6 diverse long document tasks with attribution, and experiments with different approaches to attribution on 5 LLMs of different sizes.
We find that citation, i.e. response generation and evidence extraction in one step, performs best for large and fine-tuned models, while additional retrieval can help for small, prompted models. We investigate whether the “Lost in the Middle” phenomenon exists for attribution, but do not find this. We also find that evidence quality can predict response quality on datasets with simple responses, but not so for complex responses, as models struggle with providing evidence for complex claims. We release code and data for further investigation [^1] Github repository, code under Apache 2.0, dataset licenses depend on original license, see §B..
Attribute or Abstain: Large Language Models as Long Document Assistants
Jan Buchmann, Xiao Liu, Iryna Gurevych Ubiquitous Knowledge Processing (UKP) Lab Department of Computer Science and Hessian Center for AI (hessian.AI) Technical University of Darmstadt www.ukp.tu-darmstadt.de
1 Introduction
Figure 1: Long document assistants should attribute, i.e. provide responses with evidence, or abstain. Example from QASPER Dasigi et al. (2021). Figure requires emojis to display correctly.
Figure 2: The approaches to attribution in long document scenarios analyzed in this work.
Recent LLMs can process long documents Shaham et al. (2023); Li et al. (2023b), showing great potential as long document assistants. For example (Fig. 1), such an assistant could answer a researcher’s questions about a paper. However, due to LLM hallucinations Slobodkin et al. (2023), the researcher must verify responses, which is difficult with lengthy papers. To improve verifiability and trust, the assistant should either attribute Rashkin et al. (2023) or abstain Slobodkin et al. (2023): If it finds the necessary information, it should provide a response and point to the evidence in the paper (attribute). If not, it should clearly communicate this (abstain). We investigate the capabilities of LLMs to fulfill these requirements, and the relation between response quality (i.e. correctness) and evidence quality (i.e. the relevance of the evidence to the response).
Thus far, attribution has only been investigated in retrieval augmented generation (RAG) settings, where evidence comes from a large corpus (e.g. Wikipedia) that does not fit the LLM context. This means that some form of retrieval is required to provide the LLM with a limited number of relevant passages or to retrieve evidence post-hoc. Performance depends on retrieval quality, and the best-performing approach is unclear Bohnet et al. (2022); Gao et al. (2023b); Malaviya et al. (2023).
In contrast, when the potential evidence is a single long document that fits the LLM context window, the confounding retrieval can be omitted, as LLMs can cite from their input Gao et al. (2023b). Still, it is possible that separating response generation and evidence retrieval improves performance, as shown in related works on task decomposition Sun et al. (2023) and reducing long LLM inputs, where Agrawal et al. (2024) found positive effects, while Xu et al. (2024) and Bai et al. (2023) did not. These works did not consider attribution, so there is a lack of knowledge on the effect of separating response generation and evidence retrieval, and the optimal approach to attribution on long documents.
The document length poses additional challenges: Recent works have found that LLM performance on long-input tasks depends on the position of the information in the context Liu et al. (2024); Staniszewski et al. (2023); Ravaut et al. (2023). Whether this can also be observed for attribution has not yet been investigated.
Evidence quality can be automatically evaluated without external reference Honovich et al. (2022); Yue et al. (2023); Tang et al. (2024). If evidence quality were positively correlated with response quality, bad responses could be filtered by identifying responses with bad evidence quality, thereby improving abstaining. If not, this would lead to abstaining from potentially helpful responses with insufficient evidence. Current research on the relation of response quality and evidence quality is inconclusive: Bohnet et al. (2022) and Gao et al. (2023b) reranked multiple sampled responses by evidence quality. While the former found an improvement in response quality, the latter did not. Neither provide an analysis, so we lack understanding if and how evidence quality correlates with response quality.
To close these gaps, we compile LAB, a Long-document Attribution Benchmark of 6 long-document datasets with diverse tasks (QA, classification, fact checking, NLI and summarization) and domains (science, law, governmental and Wikipedia). We conduct experiments using approaches to attribution with and without additional retrieval on 5 LLMs of varying sizes, prompted and fine-tuned, to answer three research questions:
RQ1: What are optimal approaches for attribution in long document tasks? We find that large and fine-tuned models reach best evidence quality by directly citing from input documents, but small prompted LLMs can benefit from post-hoc evidence retrieval.
RQ2: Do LLMs exhibit positional biases in evidence attribution? Concerning evidence retrieval, except for GovReport, we find no particular bias, as the predicted and gold evidence distributions are mostly similar. However, we find that response quality generally decreases as evidence appears later in the document.
RQ3: What is the relation between evidence quality and response quality? We find that evidence quality can predict response quality on datasets with single-fact responses, but not so for multi-fact responses, as models struggle with providing evidence for complex claims.
2 Related Work
Attribution
Current research in attribution is done in three strains: First, some works evaluate a range of approaches in their ability to produce attributed responses Bohnet et al. (2022); Liu et al. (2023), some proposing new datasets Malaviya et al. (2023) or benchmarks DeYoung et al. (2020a); Gao et al. (2023b). Second, methodological works propose new fine-tuning Schimanski et al. (2024); Huang et al. (2024) and prompting Berchansky et al. (2024); Fierro et al. (2024) methods to improve the citation capabilities of language models, but do not compare to approaches with additional retrieval. Both of these strains have focused on open domain QA, and neglected the long document scenario. Here, we close these gaps by providing a comprehensive investigation of attribution for long documents.
To evaluate evidence quality automatically, most works have used TRUE, Flan-T5-XXL fine-tuned on several NLI datasets Honovich et al. (2022); Bohnet et al. (2022); Gao et al. (2023b); Fierro et al. (2024); Huang et al. (2024). More recently, Attrscore Yue et al. (2023) and Minicheck Tang et al. (2024) were proposed specifically for the evaluation of attributability. We compare these models and employ the best-performing for attributability evaluation.
LLMs for long documents
While there is no universal definition of "long documents", existing long document benchmarks contain documents of 1500 to 50000 words average length Shaham et al. (2023); Dong et al. (2024); An et al. (2024); Li et al. (2023b). Initial LLMs were limited to contexts of less than 2000 tokens Brown et al. (2020); Touvron et al. (2023), but recent advances in hardware and efficiency Dao (2023) have spurred the development of models with context $\geq$8000800080008000 tokens, e.g. Longchat Li et al. (2023a), Mistral Jiang et al. (2023), GPT-3.5-16K  [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo or GPT-4-Turbo-128K. [^3] https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 We add to this line of research by evaluating a range of models in their attribution capabilities.
3 Formalization
We first define the task of producing attributed responses. Based on this, we define the approaches to attribution compared in this work.
Task definition
We assume an instruction $I$ (e.g. a question) and a document $D$ consisting of segments $d$ (e.g. paragraphs). There are two subtasks that can be solved jointly or independently: One is to generate a response $R$ (e.g. an answer) containing statements $r$. If the response is not abstained (e.g. by saying “unanswerable”), the other subtask is to retrieve evidence $E_{i}\subset D$ for each $r_{i}$ such that $r_{i}$, is attributable to $E_{i}$, i.e. "according to $E_{i}$, $r_{i}$" is true Rashkin et al. (2023).
Approaches to attribution
Different approaches to attribution can be defined based on the subtask order (Fig. 2): (1) post-hoc: an LLM generates a response $R$, and evidence is retrieved from $D$ based on $R$. (2) retrieve-then-read: Evidence $E$ is retrieved from $D$, and an LLM generates a response based on $E$. (3) citation: Based on $D$, an LLM generates a response and retrieves evidence in one step.
To decrease the input length in post-hoc and citation, $D$ can be reduced to $D^{\prime}\subset D$ via an additional initial retrieval step. This results in two further attribution approaches: (4) reduced-post-hoc and (5) reduced-citation.
4 Methods
We first introduce the datasets (§4.1) and evaluation metrics (§4.2) used in LAB, our long document attribution benchmark. We then describe our experimental setup for generation (§4.3) and retrieval (§4.4).
4.1 Datasets
Domain Task #Inst Train/Dev/Test Doc #W Res $|R|$/#W #Evi Evi Lvl
QASPER (QSP) [1] Science QA 2675/1005/1451 3937 1/12 1.7 para
Natural Questions (NQ) [2] Wiki QA 232191/6205/7307 4693 1/3 1 para
Evidence Inference (EI) [3] Science Cls 18545/1232/1218 3962 1/1 1.1 para
Wice (WIC) [4] Web FC 4234/349/358 1339 1/1 3.7 sent
ContractNLI (CNLI) [5] Legal NLI 7191/1037/2091 1697 1/1 1.5 para
GovReport (GR) [6] US Gov. Sum 15107/964/969 8464 20/517 N/A N/A
Table 1: The datasets in LAB span multiple domains and task types. The numbers for Natural Questions differ from the original publication, as we filtered the instances (§B). Column names: Doc #W: Average number of words per document. Res $|R|$/#W: Number of statements / Average number of words per response. #Evi: Number of annotated evidence segments per instance. Evi Lvl: Level of annotated evidence. Tasks: QA: Question answering. Cls: Classification. FC: Fact checking. NLI: Natural Language Inference. Sum: Summarization. [1] Dasigi et al. (2021) [2] Kwiatkowski et al. (2019) [3] DeYoung et al. (2020b) [4] Kamoi et al. (2023) [5] Koreeda and Manning (2021) [6] Huang et al. (2021)
The datasets in LAB are shown in Table 1. All datasets are in English. GovReport Huang et al. (2021) is the only dataset without annotated gold evidence. To simulate gold evidence, we use BM25 to find the 2 best-matching paragraphs from a document for each sentence in the gold summary similar to Ravaut et al. (2023). Due to limited resources, we use at most 2000 test instances from any dataset (100 for GPT-4). For details and examples see §B.
4.2 Evaluation
We evaluate two main properties of attributed responses: 1. Response quality, i.e. whether a model-generated response is correct in reference to an annotated ground-truth. 2. Evidence quality, i.e. whether the provided evidence is relevant to the given response. These are evaluated with task-specific metrics (see below). For QA datasets, we additionally evaluate Unanswerable F1, i.e. whether models abstain from responding to unanswerable instances.
4.2.1 Response Quality
For comparability with related work, we used established metrics to evaluate response quality: Exact match F1 (EM) [^4] Found to correlate better than BERTScore Zhang et al. (2020b) with human judgment in our preliminary experiments. on QASPER Dasigi et al. (2021); Shaham et al. (2023) and Natural Questions Kwiatkowski et al. (2019); Bohnet et al. (2022), classification macro F1 (CF1) for Evidence Inference DeYoung et al. (2020b), Wice Kamoi et al. (2023) and ContractNLI Koreeda and Manning (2021) and ROUGE-L [^5] ROUGE has shown good correlation with human relevance judgments in Wu et al. (2024). (RL, Lin 2004) for GovReport Huang et al. (2021); Shaham et al. (2023).
4.2.2 Evidence Quality
Evidence F1 (EF1)
For datasets that define a fixed response vocabulary (i.e. Evidence Inference, Wice and ContractNLI), we compute evidence quality as evidence F1, comparing the predicted evidence with annotated ground truth evidence. If there is no annotated evidence, evidence F1 is 1 if no evidence was predicted and otherwise 0.
Attributability (ATT)
For the other datasets, evidence F1 is insufficient: GovReport does not come with annotated evidence, and for datasets with free-form responses (QASPER and Natural Questions), evidence F1 is too rigid: A model might produce a response different from the ground truth, but supported by the retrieved evidence. Even though the predicted evidence is relevant to the predicted response, evidence F1 might be low. For these datasets, we evaluate evidence quality as attributability [^6] Also known as citation recall Gao et al. (2023b). Rashkin et al. (2023), Gao et al. (2023b); Huang et al. (2021); Schimanski et al. (2024). We assume an attributability evaluation model $\mathrm{M_{a}}(E,r)\rightarrow\{1,0\}$. The attributability score (AS) of a response is computed as the proportion of attributable statements.
$\mathrm{ATT}=\frac{1}{n}\sum_{i=1}^{n}\mathrm{M_{a}}(E_{i},r_{i})$
On QA tasks, models can abstain from responding. In these cases, we do not evaluate attributability, as no evidence is required. Because this could encourage abstaining from responding too often, we additionally evaluate Unanswerable F1 (see below).
4.2.3 Attributability Evaluation Model Selection
To select a model for attributability evaluation, we created test datasets for QASPER, Natural Questions and GovReport, and evaluated TRUE Honovich et al. (2022), AttrScore Yue et al. (2023) and Minicheck Tang et al. (2024). For AttrScore, we map "Contradictory" and "Extrapolatory" predictions to a single "not attributable" class.
Human annotation
We generated attributed responses using GPT 3.5-post-hoc with BM25 for evidence retrieval. Two authors of this study, both holders of a Master’s degree and fluent in English, annotated attributability (attributable or not attributable) for 200 responses (200 sentences for GovReport) and reached an agreement of 0.74 (QASPER), 0.77 (Natural Questions) and 0.76 (Govreport) Cohen’s $\kappa$.
Results
We report accuracy Gao et al. (2023b) and balanced accuracy Tang et al. (2024) scores in Table 2. The scores are comparable to Gao et al. (2023b), who reported 85% accuracy and Tang et al. (2024), who reported between 59% and 84% balanced accuracy on their respective benchmarks. We select the best-performing model to evaluate attributability: Minicheck for QASPER, and TRUE for Natural Questions and GovReport.
4.2.4 Unanswerable F1 (UF1)
We set this up as a classification task similar to Slobodkin et al. (2023). Gold labels are determined depending on the number of annotations Kwiatkowski et al. (2019): For 3 annotations or less: “Unanswerable” if all annotators annotated unanswerable, else “answerable”. More than 3 annotations: “Unanswerable” if at most one annotator did not annotate unanswerable, else “answerable”. If a model abstains, its prediction is set to “unanswerable”, else “answerable”.
To detect abstaining, we compiled a list of keyphrases based on Slobodkin et al. (2023) and manual inspection. If a response contained a keyphrase, any predicted evidence was removed, and the response was set to "unanswerable" (see §G).
Model QSP NQ GR
TRUE 79/80 83/83 79/78
AttrScore 76/71 68/67 76/52
Minicheck 83/82 82/82 79/70
Table 2: Attributability evaluation model selection. Metrics: accuracy / balanced accuracy Tang et al. (2024)
4.3 Generation
Model selection
We focus on two groups of LLMs with at least 8K tokens input length: (1) The large state of the art models GPT-3.5777gpt-35-turbo-0613-16k and GPT-4888gpt-4-turbo-128k, as they hold top positions in other long document benchmarks Li et al. (2023b); Shaham et al. (2023); An et al. (2024) (2) Small (~3-7B) models that are accessible with limited resources. We tested a range of models with citation on QASPER and GovReport and selected Longchat [^9] Longchat-7B-v1.5-32K Li et al. (2023a) and Mistral [^10] Mistral-7B-Instruct-v0.2 Jiang et al. (2023) as the best-performing for prompting and Flan-T5  [^11] Flan-T5-XL Longpre et al. (2023) as the best performing for fine-tuning. For complete selection results and hyperparameters see §E.
Prompts
We employ separate prompt sets for citation and non-citation. Similar to Shaham et al. (2023), we keep instructions short, including guidance on the expected responses and output format. Prompts contained three in-context examples, where documents were shortened to title, section headings and annotated evidence. For details and prompt optimization experiments, see §C.
Response parsing
For all datasets except GovReport, responses consist of single statements. For GovReport, we split responses into statements (sentences) using NLTK Loper and Bird (2002). For citation, LLMs are expected to generate segment identifiers (“[1] [2]”) at the end of each statement.
4.4 Retrieval
Retriever selection
For retrieval in post-hoc, retrieve-then-read and reduced, we employ sparse and dense retrievers that showed good performance in related work [^12] For efficiency reasons, we do not use LLMs for retrieval. Thakur et al. (2021): BM25 Robertson and Zaragoza (2009); Gao et al. (2023b), GTR Ni et al. (2022); Gao et al. (2023b); Bohnet et al. (2022), Contriever Izacard et al. (2021); Xu et al. (2024); Bai et al. (2023), Dragon Lin et al. (2023); Xu et al. (2024) and the best-performing Sentence Transformer “all-mpnet-base-v2” [^13] https://sbert.net/docs/sentence_transformer/pretrained_models.html#semantic-search-models Reimers and Gurevych (2019). For each combination of approach and task, we selected the best-performing retriever using GPT-3.5 as response generator (see §F).
Details
In retrieve-then-read and reduced, queries were constructed based on information available in the instruction (e.g. question or a claim). For GovReport, similar to Zhang et al. (2020a), we created queries from all document segments and retrieved paragraphs based on self-similarity. In post-hoc, queries were constructed based on the instruction and the generated response. For both post-hoc and retrieve-then-read, we retrieve 5 evidence segments for Wice, and 2 for all other datasets based on evidence statistics (see Table 1). In reduced approaches we reduce input documents to 10 segments based on Xu et al. (2024) (see §F).
5 Experiments
QSP NQ EI WIC CNLI GR Avg EM ATT UF1 EM ATT UF1 CF1 EF1 CF1 EF1 CF1 EF1 RL ATT RQ EQ GPT-3.5 p-h 45 62 65 51 44 58 77 22 52 48 44 40 26 74 49.17 48.33 rtr 42 78 51 50 42 57 73 25 50 44 47 41 21 40 47.17 45.00 cit 51 55 70 47 38 53 78 50 52 59 43 53 26 59 49.50 52.33 r-p-h 48 71 58 50 45 57 78 22 52 48 45 40 22 64 49.17 48.33 r-cit 50 69 64 46 38 53 78 52 48 63 48 57 22 58 48.67 56.17 GPT-4 p-h 65 68 68 52 42 56 87 24 32 36 66 50 27 63 54.83 47.17 rtr 56 83 56 58 36 61 74 31 22 28 64 59 20 27 49.00 44.00 cit 68 76 71 51 49 57 86 64 35 47 63 64 27 62 55.00 60.33 r-p-h 62 71 69 51 43 57 84 24 28 30 65 50 22 54 52.00 45.33 r-cit 63 76 66 54 49 58 83 49 33 46 64 67 22 55 53.17 57.00 Flan-T5 p-h 55 67 58 81 61 74 86 22 44 46 77 57 27 90 61.67 57.17 rtr 43 74 55 79 62 74 77 25 43 43 64 58 19 67 54.17 54.83 cit 53 57 56 84 75 71 83 59 53 70 77 78 25 71 62.50 68.33 r-p-h 53 72 58 80 62 75 87 22 45 46 76 58 23 84 60.67 57.33 r-cit 52 62 55 84 75 73 84 57 53 68 75 75 22 68 61.67 67.50 Longchat p-h 23 57 62 21 33 38 66 22 23 31 38 33 24 67 32.50 40.50 rtr 29 56 49 27 32 43 43 25 19 25 39 39 21 41 29.67 36.33 cit 21 6 60 17 2 38 66 13 24 19 36 14 25 9 31.50 10.50 r-p-h 33 58 54 25 35 41 61 22 20 27 33 30 22 63 32.33 39.17 r-cit 26 26 53 21 4 42 65 13 20 25 24 12 22 23 29.67 17.17 Mistral p-h 32 67 59 28 42 43 75 20 37 45 41 40 22 63 39.17 46.17 rtr 27 76 50 30 42 46 66 25 37 42 51 43 20 39 38.50 44.50 cit 35 39 61 26 15 41 75 27 34 39 47 38 22 0 39.83 26.33 r-p-h 31 71 58 29 44 43 76 21 38 45 45 38 21 64 40.00 47.17 r-cit 36 53 58 28 21 42 77 26 35 47 51 44 22 1 41.50 32.00
Table 3: Evaluation on LAB, all scores show percentages. Citation / reduced-citation mostly perform best, with notable exceptions for GovReport and Longchat (see §5.1). p-h: post-hoc. rtr: retrieve-then-read. cit: citation. r-: reduced-. EM: Exact match F1, ATT: Attributability, UF1: Unanswerable F1, CF1: Classification F1, EF1: Evidence F1, RL: Rouge-L, RQ (EQ): Average response (evidence) quality, mean of all scores with blue (green) shade.
5.1 RQ1: What are optimal approaches to attribution in long document tasks?
Table 3 shows the results from all combinations of selected models. Due to the large number of experiments, all results are from single runs.
Which approach produces the highest evidence quality?
Flan-T5-XL has higher average scores than GPT-3.5 and GPT-4, while the Longchat and Mistral scores are lower. For GPT-3.5, GPT-4 and Flan-T5, citation / reduced-citation results in the best evidence quality on average and most datasets, and retrieve-then-read performs worst. Post-hoc works best for Longchat and Mistral, and for all models on GovReport.
Does citation hurt response quality?
It could be assumed that post-hoc results in better response quality than citation, as task decomposition can improve performance Gao et al. (2023a). We compare average response quality between (reduced-)citation and (reduced-)post-hoc for GPT-3.5, GPT-4 and Flan-T5-XL. In no case, the response quality for citation is more than 0.5 points lower than for post-hoc, showing that citation has a minimal effect on response quality.
Does performance depend on document length?
Tables 7 and 8 show the correlation between response quality or evidence quality and document length, which is mostly negative. A notable exception is GPT-4 evaluated for response quality, where correlation with document length is positive on all tasks but GovReport.
Does reduction of the input document help?
Comparing reduced-post-hoc to post-hoc and reduced-citation to citation, we find that response quality is mostly better for the non-reduced variant, with the exception of Mistral. Regarding evidence quality, reduction only helps for GPT-3.5-citation, Longchat-citation and -post-hoc and Mistral-citation.
Discussion
Citation or reduced-citation result in the best average evidence quality, while not hurting response quality, in line with recent work showing LLM capabilities for retrieval Ma et al. (2023). The GovReport task and the small models Longchat and Mistral are exceptions to this, as post-hoc results in better evidence quality in these cases. For GovReport, the higher evidence quality with post-hoc can be explained with the "repetitive" nature of the summarization task, since the high overlap between response statements and document provides good conditions for retrievers to find evidence. For the small models, related work has shown that they lack instruction following capability to perform evidence extraction Gao et al. (2023b); Schimanski et al. (2024), making post-hoc the better approach for the model.
Comparing models, fine-tuned Flan-T5-XL has higher average scores for response and evidence quality than the large prompted models GPT-3.5 and GPT-4. This could also be observed in related work Huang et al. (2024); Schimanski et al. (2024).
Regarding the relation between input length and performance, we mostly found negative correlation, similar to Bai et al. (2023) and Kwan et al. (2023). Still, similar to Xu et al. (2024) and Bai et al. (2023), we have not found a general beneficial effect of input reduction on performance. The positive results from Agrawal et al. (2024) were obtained in a multi-document setting, where there is no logical coherence between the input documents. In contrast, the logical coherence in long documents can be disrupted through reduction, which can make processing of the reduced document more difficult.
5.2 RQ2: Do LLMs exhibit positional biases in attribution?
Figure 3: Top: Evidence distribution (predicted via citation) by position in the document. Except for GovReport, no positional bias is visible. Bottom: Response quality by position of gold evidence in the document. Negative correlation between evidence position and response quality is visible in several cases (Table 5) (see §5.2).
Several works have shown that LLM performance depends on the position of information in the input Liu et al. (2024); Staniszewski et al. (2023); Ravaut et al. (2023). We investigate whether this phenomenon exists for attribution.
Do predicted and gold evidence distributions agree?
Figure 3 (top) shows the predicted evidence distributions from citation [^14] We focus on citation because only in this approach the LLM performs evidence retrieval. and the gold distributions. [^15] For GovReport, we matched summary sentences to paragraphs using BM25 to obtain gold evidence, see §4.1 The models generally follow the gold evidence distribution, with Longchat showing the strongest deviation. GovReport is an exception: all models show a higher focus on the beginning of the document, especially Flan-T5-XL. Figures 5 and 6 show the relation of evidence quality and the position of gold or predicted evidence, respectively. Neither of them show a clear dependence of evidence quality on evidence position.
Does response quality depend on the position of gold evidence?
We grouped instances by evidence position and evaluated the approaches with full document input (citation and post-hoc) separately for each group (Fig. 3, bottom). The strong fluctuations of GPT-4 can be explained by the fact that it was evaluated on 100 samples only. It seems that response quality decreases as evidence appears later in the document. Similar to Ravaut et al. (2023), we computed Spearman correlation between response quality and evidence position, and find that correlation is mostly negative, but not always significant (Table 5).
Discussion
Except for GovReport, we could not find positional biases in LLM evidence retrieval. From the results of Liu et al. (2024), we expected to find a “Lost in the Middle” effect, i.e. a reduction of retrieved evidence or performance in the middle of documents. Rather, we found a decrease in response quality towards the end of the document similar to Ravaut et al. (2023). We and Ravaut et al. (2023) work with coherent long documents, while Liu et al. (2024) work in a RAG-like setup, where the input are multiple documents without coherence, and models might expect an ordering by relevance. This might explain the different results.
5.3 RQ3: What is the relation between evidence quality and response quality?
QSP NQ EI WIC CNLI GR
EM UF1 EM UF1 CF1 CF1 CF1 RL
GPT-3.5 post-hoc 3 2 17 20 12 4 19 0
citation 2 1 14 18 13 4 25 -1
GPT-4 post-hoc 4 1 20 27 7 10 16 0
citation 4 3 6 18 10 7 20 1
Flan-T5 post-hoc 0 2 13 12 7 1 8 0
citation 2 1 11 10 11 3 8 -1
Longchat post-hoc 6 2 13 20 18 16 26 0
citation 4 1 4 7 10 11 15 1
Mistral post-hoc 6 2 13 19 9 14 23 1
citation 6 1 9 12 9 12 15 0
Table 4: Difference in response quality-coverage AUC between responses ordered by evidence quality (attributability) and random ordering. Evidence quality predicts response quality on several datasets (see §5.3)
Figure 4: Gold class distribution in selective prediction for GPT-3.5-citation.
Attributability evaluation models Honovich et al. (2022); Yue et al. (2023); Tang et al. (2024) can evaluate evidence quality without external reference. If evidence quality were positively correlated with response quality, this could be used to abstain from low-quality responses. We test this with selective prediction El-Yaniv and Wiener (2010), using evidence quality scores from attributability evaluation as an estimate of confidence.
Selective Prediction
We begin by filtering predicted responses that don’t require evidence161616“Unanswerable” on QASPER and Natural Questions, “not supported” on Wice, “not mentioned” on ContractNLI., as attributability cannot be evaluated on these. We sort the remaining $l$ predictions by evidence quality [^17] We evaluated evidence quality as attributability for all datasets, as it does not require annotated evidence. We used MinicheckTang et al. (2024) for QASPER and TRUE Honovich et al. (2022) for all other datasets (§4.2.3) in descending order, obtaining the ordered set of responses $\mathcal{R}_{sel}$. For each $t\in\{0...l\}$, we compute average response quality [^18] As filtering responses that do not require evidence produces a strong class imbalance, we use micro F1 instead of macro F1 for Wice and ContractNLI on the subset of $\mathcal{R}_{sel}$ up to $t$ and the coverage (the proportion of responses evaluated). We evaluate the estimation of confidence through evidence quality by computing the area under the response quality-coverage curve (AUC) Chen et al. (2023).
Is evidence quality a good estimate of confidence in selective prediction?
Table 4 shows the AUC difference between ordering predictions by attributability and a random order baseline (mean of 10 repeats). We see that attributability is an effective estimate of confidence on Natural Questions, Evidence Inference and ContractNLI, and to some extent on Wice. On QASPER, and GovReport, the difference to the random baseline is small.
Do unanswerable instances have lower evidence quality?
For instances annotated as “unanswerable”, “not supported” or “not mentioned”, there is no annotated evidence. If models give different responses, the evidence quality should be low, and these should be filtered in selective prediction. Fig. 4 shows the distribution of gold responses in selective prediction. As expected, the proportion of such instances without sufficient evidence decreases with lower coverage for Natural Questions, Wice and ContractNLI but, surprisingly, not for QASPER.
Why does evidence quality fail to predict response quality?
We consider GovReport a special case, as its long responses are evaluated in their entirety, which might be too coarse-grained to reflect the per-statement evaluation of evidence quality. This is corroborated by the fact that system-level correlation between evidence quality and response quality is significantly positive for all datasets except GovReport (Table 6).
For Wice and QASPER, the possible causes are: (1) Responses are correct, but the evidence is insufficient. (2) Responses are incorrect, but the evidence is sufficient. (3) The attributability scores are wrong. We performed a manual analysis on the 50 responses with the lowest attributability from GPT-3.5-citation: For QASPER, the responses were often correct (answer F1 of 66), but the evidence was insufficient in 46 cases. For Wice, this could be observed in 39 cases. This implies that the LLM’s failure to extract sufficient evidence is the main reason for low correlation between evidence and response quality.
For Wice, the attribution evaluation model failed to recognize evidence for “partially supported” claims in 8 cases, as it is only trained to distinguish “supported” and “not supported”. This can be seen in Fig. 4, where the proportion of “partially supported” decreases with lower coverage.
Discussion
We explain the differences in the relationship between evidence quality and response quality by the varying dataset complexity. While Natural Questions, ContractNLI and Evidence Inference focus on single facts (e.g. a single entity or a specific contractual obligation), Wice and QASPER instances often contain multiple facts (e.g. an enumeration or multiple subclaims), which is also reflected in the higher number of evidence segments per instance (Table 1). Models respond correctly, but fail to point to all necessary evidence. This is in line with related work on attribution: While Bohnet et al. (2022) found that response quality can be improved by attributability-based reranking on Natural Questions, Gao et al. (2023b) did not find this on their more complex benchmark.
6 Conclusion
In our experiments on LAB, we found that citation is a promising approach to attribution for large or fine-tuned models, while for small prompted models, post-hoc extraction can improve performance. We did not find a “Lost in the Middle” effect, but negative correlation between evidence position and response quality in some cases. Finally, we showed that evidence quality can predict response quality for responses with low complexity. We hope that our results, code and data spur further research on long document assistants, most prominently: (1) Improving the citation capabilities of LLMs for complex responses. (2) Combining attributability evaluation models and iterative self-refinement Gao et al. (2023a) to try to improve abstained responses.
7 Limitations
Using LLMs for retrieval
LLMs have shown good performance in reranking tasks (e.g. Ma et al. 2023). For efficiency reasons, we did not employ LLMs for retrieval. Instead, we employed state-of-the-art retrievers and implemented a rigorous selection procedure, elucidating the best retriever for each combination of task and approach. In the case of reduced approaches, we deem it unlikely to see large beneficial effects: The “pressure” on the retrievers was already low, as they only had to retrieve 1-3 relevant segments among 10 retrieved in total.
For post-hoc and retrieve-then-read, it could be that using LLMs for evidence retrieval improves performance. However this would not change our main claims: Our experiments with the citation approach already show that using LLMs for evidence retrieval works best. Therefore, we found the trade-off between increased computational cost and additional insights not favorable towards employing LLMs for evidence retrieval in post-hoc and retrieve-then-read. This might be different for researchers or practicioners interested in maximal performance.
Evaluation of evidence quality
While our attributability model selection experiments showed that these models obtain good accuracy, our analysis in 5.3 showed that edge cases are not yet handled well. Research into solving such edge cases is a promising direction for future work.
Datasets
We compiled a benchmark with diverse tasks, domains, response and document lengths, but naturally, we were not able to cover all variations of these properties. Many long document datasets and benchmarks are available (e.g. Li et al. 2023b; An et al. 2024; Shaham et al. 2023), but only few contain annotated evidence, which we required for the positional bias analysis in our paper (not finding a long document summarization dataset with annotated evidence, we resorted to GovReport). Given that we found only limited positional biases, extending our work to more datasets with and without annotated evidence is an interesting direction for future work.
8 Potential Risks
The goal of this work is to promote the development of more trustworthy long document assistants, which we deem a promising, but low-risk research goal. Similarly, we deem our research methodology to be of low risk: All datasets used were created for NLP research, do not contain personal, harmful or sensitive data and were published under permissive licenses (Table 9). We did not employ human annotators other than the authors of the study themselves.
9 Acknowledgements
This work was funded by the European Union (ERC, InterText, 101054961). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work has been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
Appendix A Additional Results
A.1 Correlation between Evidence Position and Response Quality
See Table 5.
QSP NQ EI WIC CNLI GR
GPT-3.5 post-hoc -0.76* -0.03 -0.77* -0.49 0.42 -0.45
citation -0.08 -0.44 -0.62 -0.41 -0.02 -0.59
GPT-4 post-hoc 0.12 0.02 0.03 0.35 0.04 0.12
citation 0.12 0.04 0.03 0.09 -0.19 0.03
Flan-T5 post-hoc -0.32 0.05 -0.75* -0.37 -0.07 -0.92*
citation -0.19 -0.13 -0.81* -0.76* -0.26 -0.95*
Longchat post-hoc -0.43 0.44 -0.33 -0.32 -0.58 -0.64*
citation -0.49 -0.28 -0.42 -0.30 -0.54 -0.61
Mistral post-hoc -0.53 0.39 -0.59 0.04 0.09 -0.92*
citation -0.35 0.31 -0.37 -0.58 0.47 -0.82*
Table 5: Pearson correlation between response quality and position of annotated evidence. * indicates significance ($p<0.05$). See §5.2 for details.
A.2 System-level Correlation of Attributability and Response Quality
Pearson $r$ ($p$)
QSP $0.93$ ($7.7\times 10^{-5}$)
NQ $0.99$ ($3.6\times 10^{-8}$)
EI $0.93$ ($7.9\times 10^{-5}$)
WIC $0.96$ ($1.2\times 10^{-5}$)
CNLI $0.93$ ($1.2\times 10^{-4}$)
GR $0.36$ ($3.0\times 10^{-1}$)
mean $0.98$ ($1.3\times 10^{-6}$)
Table 6: System-level Pearson correlation between response quality and evidence quality for citation and reduced-citation (8 score pairs per dataset). For Avg, correlation was computed over the average scores (rightmost columns in Table 3).
See Table 6
A.3 Overlap between citation and post-hoc evidence
A.4 Relation between Performance and Input Document Length
See Tables 7 and 8. See §5.1 for analysis and discussion.
QSP NQ EI WIC CNLI GR
GPT-3.5 citation 0.04 -0.01 -0.08* -0.04 0.05* -0.08*
post-hoc 0.08* -0.00 -0.05 -0.05 0.05* -0.04
GPT-4 citation 0.29* 0.11 0.13 0.08 0.34* -0.34*
post-hoc 0.28* 0.17 0.13 0.11 0.39* -0.26*
Flan-T5 citation -0.02 -0.10* -0.07* 0.05 0.00 -0.07*
post-hoc -0.03 -0.10* -0.05 -0.08 -0.02 -0.11*
Longchat citation -0.05 -0.05* -0.04 0.01 0.04 -0.15*
post-hoc -0.02 -0.13* -0.04 0.03 -0.06* -0.13*
Mistral citation -0.03 -0.07* -0.07* -0.07 0.00 -0.12*
post-hoc -0.01 -0.06* -0.07* -0.04 -0.02 -0.06
Table 7: Pearson correlation between response quality and document length. * indicates significance ($p<0.05$). See §5.1 for analysis and discussion.
QSP NQ EI WIC CNLI GR
GPT-3.5 citation -0.08* -0.13* -0.10* -0.20* -0.04 -0.24*
post-hoc -0.03 -0.10* -0.01 -0.11* -0.07* -0.23*
GPT-4 citation 0.04 -0.17 -0.09 0.09 0.15 -0.10
post-hoc -0.13 -0.05 -0.07 0.10 0.01 -0.12
Flan-T5 citation -0.04 -0.08 -0.10* -0.09 -0.07* -0.11*
post-hoc -0.06* -0.20* -0.00 -0.15* -0.13* -0.10*
Longchat citation -0.13* -0.09* -0.12* -0.05 0.08* -0.03
post-hoc -0.01 -0.13* 0.00 0.03 0.04* -0.18*
Mistral citation -0.18* -0.17* -0.13* -0.19* -0.09* -0.05
post-hoc -0.08* -0.10* -0.04 -0.16* -0.06* -0.10*
Table 8: Pearson correlation between evidence quality and document length. * indicates significance ($p<0.05$). See §5.1 for analysis and discussion.
A.5 Effect of Evidence Position on Evidence Quality
To analyze the effect of the position of gold or predicted evidence on evidence quality, we grouped predictions from models under the citation approach into 5 bins based on the relative position of gold or predicted evidence in the document. We then computed mean evidence quality separately for each bin. Figures 5 and 6 show the results. A “Lost in the Middle” effect is not visible.
Figure 5: Evidence quality by position of gold evidence in the document. “Lost in the Middle” effect is not visible. For more information, see §A.5.
Figure 6: Evidence quality by position of predicted evidence in the document. A “Lost in the Middle” effect is not visible. For more information, see §A.5.
Appendix B Datasets
QASPER CC-BY 4.0
Natural Questions CC-BY-SA 3.0
Evidence Inference MIT
Wice CC-BY-SA 4.0 (text), ODC-BY (annotations)
ContractNLI CC-BY 4.0
GovReport No copyright
Table 9: Licenses of the datasets in LAB.
See Table 10 for examples from the datasets in LAB
QASPER
Dasigi et al. (2021) is a dataset of NLP papers and questions about them. Answers can be extractive, abstractive, “Yes”, “No” or “unanswerable”. Evidence is annotated on paragraph level. We remove instances with evidence in tables or figures.
Natural Questions
Kwiatkowski et al. (2019) is a dataset of genuine questions from Google search logs and Wikipedia pages that may or may not contain the answers. We removed all annotations with answers in tables and those that only have a long answer, keeping only the annotations that have short answers (i.e. extractive spans), “Yes”/“No” answers or “unanswerable”. All non-unanswerable annotations have a single evidence paragraph. As the official test set is hidden, we used the dev set for testing and a part of the train set for development.
Evidence Inference
DeYoung et al. (2020b) consists of reports from clinical studies, "prompts" in the form of intervention, comparator, and outcome, one or multiple labels for the prompt ("significantly increased", "significantly decreased", or "no significant difference") and corresponding evidence spans. We map the annotated evidence spans to paragraphs.
Wice
Kamoi et al. (2023) is a dataset of claims from Wikipedia and referenced webpages. Claims are annotated as “supported”, “partially supported” or “not supported”. The referenced webpages are annotated with evidence on sentence level. We use the full-claim subset.
ContractNLI
Koreeda and Manning (2021) is a dataset of non-disclosure agreement contracts and claims about these agreements. The relation between contract and claim is annotated as “entailment”, “contradiction” or “not mentioned”. We split the contract documents into paragraphs at newline symbols to obtain paragraphs, and map the sentence-level annotated evidence to these paragraphs.
GovReport
Huang et al. (2021) is a dataset of reports from US-American governmental institutions and their executive summaries.
Dataset format
All datasets were converted to the Intertext Graph format Kuznetsov et al. (2022) to enable shared processing and the use of document structure (where available).
Appendix C Prompts
The prompts used in our experiments can be divided into two building blocks: (1) Instructioncand (2) instance specific input. The instruction further consists of (a) task explanation and (b) format explanation. We explain the building blocks in the following. For a complete prompt example, see Table 15.
C.1 Instruction
C.1.1 Task explanation
Task explanations give information on the type of document used as input, the type of task to be solved, and possible labels. Table 11 shows the task explanations used.
C.1.2 Format explanation
Format explanations were only used for citation approaches, explaining the expected format of pointers to evidence segments. Table 12 shows the format explanations used for single-statement responses and multi-statement responses. The single statement explanation was used for all datasets except GovReport, where the multi statement explanation was used.
C.2 Instance Specific Input
Instance specific input consisted of an input document and task-dependent additional information, such as a question or a claim. Table 13 shows the formatting of instance specific input.
C.3 Example document formatting
We shortened the documents in examples to document title, section headings (where available) and annotated evidence segments. If there were no annotated evidence segments (e.g. because an example instance is unanswerable) we selected 2 random segments from the document (5 for Wice).
C.4 Prompt Selection
We optimized two prompt properties: The position of the instruction (i.e. task explanation and format explanation) and the number of few-shot examples. We ran experiments employing GPT-3.5 on QASPER and GovReport under the citation approach, the results are shown in Table 14. We first varied the position of the instruction, finding that the instruction before the instance specific input resulted in best performance. Next, we experimented with using 1, 2 or 3 few shot examples, finding that 3 examples resulted in best performance. We limited the number of few-shot examples to leave enough space for the input document.
C.5 Complete prompt example
See Table 15.
Appendix D Attributability Evaluation
To evaluate attributability, we experimented with TRUE Honovich et al. (2022), Attrscore (Flan-T5-XXL version) Yue et al. (2023) and Minicheck (Flan-T5-Large version) Tang et al. (2024). These models expect a claim and evidence as input. In the following we explain the construction of claims, evidence formatting and model-specific prompts.
D.1 Claim Construction
Claims were constructed based on the task specific inputs and outputs. Table 16 shows examples.
QA datasets
For QASPER and Natural Questions, question and answer were concatenated to get the claim.
Evidence Inference
If the response was “no significant difference”, the claim was formulated as “There was no significant difference between the effect of {intervention} and the effect of {comparator} on {outcome}.”. If the response was “significantly increased” or “significantly decreased” were predicted, the claim was formulated as “The {intervention} {response} {outcome} in comparison to {comparator}”.
Wice
If the response was “supported”, the input claim was used as the claim for attributability evaluation. If the response was “partially supported”, the claim for attributability evaluation was formulated as “The claim {claim} is partially supported.” “Not supported” responses did not require attributability evaluation.
ContractNLI
If the response was “entailment”, the input claim was used as the claim for attributability evaluation. As there are only 20 claims in the complete dataset, we formulated an inverse version of each claim. This was used as the claim when the response was “contradiction”. See Table 16 for an example.
GovReport
The generated summary sentences were used as claims.
D.2 Input Formatting
Evidence formatting
Predicted evidence segments were ordered by occurence in the document, joined by newline symbols, and prepended with the document title.
Model-specific formatting
The prompt templates for attributability evaluation can be seen in Table 17. They were taken from the respective original publications.
D.3 Annotation Instructions
As mentioned in 4.2.3, we manually annotated predictions on QASPER, Natural Questions and GovReport. The annotation instructions are shown in Fig. 7.
Appendix E Generation
E.1 Model Selection
To select open source models for prompting and fine-tuning, we compared their performance in preliminary experiments. Table 18 shows all open source models considered.
Fine-tuning
We fine-tuned all candidate models for 1000 steps on QASPER and evaluated them on the dev set. As the results in Table 19 show that Flan-T5-XL has a clear advantage over the other models, we used it in all further fine-tuning experiments.
Prompting
We evaluated all candidate models on 100 instances from the QASPER and GovReport dev sets, using the citation approach with the prompts described in §C. Table 20 shows the results. As the Longchat model obtained the highest average score, we used it in all further experiments.
E.2 Hyperparameters
Generation
We set the maximum input length to 16K, truncating the input document if needed. We performed greedy decoding and temperature 0 for best reproducibility (§G).
Fine-tuning
We employed LoRA fine-tuning Hu et al. (2022) in a citation setting and a non-citation setting for 1000 steps. We set $r=64$, $\alpha=16$, a dropout rate of $0.1$, a learning rate of $410^{-}$ 4, effective batch size of $8$, and employed an AdamW optimizer Loshchilov and Hutter (2019)
Appendix F Retrieval
F.1 Retriever Selection
Candidates
We experimented with BM25 Robertson and Zaragoza (2009), GTR Ni et al. (2022), Contriever Izacard et al. (2021), Dragon Lin et al. (2023) and the best-performing Sentence Transformer “all-mpnet-base-v2” [^19] https://sbert.net/docs/sentence_transformer/pretrained_models.html#semantic-search-models Reimers and Gurevych (2019).
Results
We tested all combinations of post-hoc (Table 23), retrieve-then-read (Table 21) and reduced-citation (Table 22) approaches, tasks and retrievers, using GPT-3.5 to generate responses. We selected the best-performing retriever for each combination
F.2 Query Construction
Post-hoc query construction
Post-hoc queries were constructed by combining instance specific inputs and outputs. The exact query construction depended on the task. For QASPER and Natural Questions, question and generated response were concatenated. For Evidence Inference, ContractNLI and GovReport, post-hoc queries were constructed in the same manner as claims for attributability evaluation (see §D). For Wice, the input claim was used as the query. See Table 24 for examples.
Reduce and retrieve-then-read query construction
Queries for Reduce and retrieve-then-read were constructed based on instance specific input. For QASPER and Natural Questions, this was the question. For Evidence Inference, this was the question formed out of intervention, comparator and outcome. For Wice and ContractNLI, this was the claim. For GovReport, these were the document paragraphs. See Table 25 for examples.
Reduce and retrieve-then-read for GovReport
To find the most relevant paragraphs from documents in the GovReport dataset, we used each paragraph as a query and computed the retrieval score for all paragraphs (including the paragraph itself), resulting in $n^{2}$ scores $s_{i,j}$ for a document with $n$ paragraphs. Each $s_{i,j}$ is the score for retrieving $p_{j}$ with query $p_{i}$. We compute a single score for each paragraph as $s_{j}^{*}=\sum_{i=0}^{n}s_{i,j}$, i.e. the sum of scores to retrieve $p_{j}$. The paragraphs with the highest $s^{*}$ are then selected.
Appendix G Technical Details
Technical setup
GPT-35 and GPT-4 were accessed via the Azure OpenAI API  [^20] https://azure.microsoft.com/en-us/products/ai-services/openai-service, all other models were downloaded and run locally via Huggingface Transformers Wolf et al. (2020) on NVIDIA A100 and H100 GPUs.
Rouge Scoring
We used the rouge-score package [^21] https://pypi.org/project/rouge-score/ to evaluate ROUGE-L
Compute Budget
We spent around \$400 to access OpenAI models, including preliminary experiments. We estimate to have spent 300 GPU hours on all experiments, including fine tuning, inference and attributability evaluation.
Use of AI assistants
We used Github Copilot [^22] https://github.com/features/copilot for coding assistance.
Detecting abstaining
Based on Slobodkin et al. (2023) and our own inspection of model responses, we compiled two sets of keyphrases to detect abstaining. If any of these key was found in a response, it was set to “unanswerable”. The first set of keyphrases is shown in Table 26, and was used as is. The second set of keyphrases was constructed by combining all verbs and adverbs in Table 27 as follows:
  1. 1.
    “not <verb>”
  2. 2.
    “not <adverb> <verb>”
Verbs were used in their base form and as participles (e.g. “mention” $\rightarrow$ “mentioned”). This means that each combination of verb and adverb results in four keyphrases. For example, the verb “provide” and the adverb “explicitly” resulted in “not provide”, “not explicitly provide”, “not provided”, “not explicitly provided”.
Input Output
QASPER Question: Which domains do they explore? news articles related to Islam and articles discussing Islam basics
Natural Questions Question: who won the 2017 ncaa mens basketball tournament North Carolina
Evidence Inference Question: With respect to Quality of life, characterize the reported difference between patients receiving good motivation/capability and those receiving inadequate motivation/capability.Choose between ’significantly decreased’, ’no significant difference’, and ’significantly increased’. no significant difference
Wice Claim: Having over 3,000 animals of nearly 400 different species, the zoo has slowly increased its visitors and now ranks as the number one outdoor tourist attraction in the state.
Additional Info: The Sedgwick County Zoo is an AZA-accredited wildlife park and major attraction in Wichita, Kansas. Founded in 1971 with the help of the Sedgwick County Zoological Society, the zoo has quickly become recognized both nationally and internationally for its support of conservation programs and successful breeding of rare and endangered species.
Partially Supported
ContractNLI Claim: Receiving Party shall not reverse engineer any objects which embody Disclosing Party’s Confidential Information. not mentioned
GovReport Question: Write a one-page summary of the document. Structure your summary according to the following questions: 1. Why GAO Did This Study 2. What GAO Found 3. What GAO Recommends {summary}
Table 10: Dataset examples
QASPER You are given a Scientific Article and a Question. Answer the Question as concisely as you can, using a single phrase or sentence. If the Question cannot be answered based on the information in the Article, answer "unanswerable". If the question is a yes/no question, your answer should be "yes", "no", or "unanswerable". Do not provide any explanation. (If the question can be answered, provide one or several evidence paragraphs that can be used to verify the answer. Give as few paragraphs as possible.)
Natural Questions You are given a Wikipedia page and a question about that page. Answer the question as concisely as you can, using at most five (5) words. If the question cannot be answered based on the information in the article, write "unanswerable". If the question is a yes/no question, answer "yes", "no", or "unanswerable". Do not provide any explanation. (If the question can be answered, provide one evidence paragraph that can be used to verify the answer.)
Evidence Inference You are given a clinical study report and a question. Assess the effect of a treatment on a clinical outcome, compared to a control treatment. The options are "significantly increased", "significantly decreased" and "no significant difference". Do not provide any explanation. (Provide one or several evidence paragraphs that can be used to verify the answer. Give as few paragraphs as possible.)
Wice You are given a document and a claim. Evaluate if the claim is supported by the document. You can choose between “supported”, “partially supported” and “not supported”. Do not add any explanation. (If you answer “supported” or “partially supported”, provide the evidence sentences from the document that can be used to verify the answer. Give as few sentences as possible.)
ContractNLI You are given a non disclosure agreement contract and a statement. Determine the relation between the contract and the statement. You can choose between “entailment”, “contradiction” and “not mentioned”. Do not add any explanation. (If you answer “entailment” or “contradiction”, provide the evidence paragraphs from the contract that can be used to verify the answer. Give as few paragraphs as possible.)
GovReport You are given a government report document. Write a one page summary of the document. (Each sentence in your summary should reference the source paragraphs from the document that can be used to verify the summary sentence.)
Table 11: Task explanations for the datasets in LAB. Text in parentheses at the end was only shown when citation approaches were used.
Single statement Your reply must have the following format:
"<answer> [X] [Y]"
In your reply, replace <answer> with your solution to the task. Your reply must be followed by the ids of the relevant segments from the document.
Multi statement Your reply must have the following format:
"<answer_sentence_1>[X] [Y] <answer_sentence_2>[Z]…"
In your reply, replace <answer_sentence_1> with your first sentence, <answer_sentence_2> with your second sentence, and so forth. Each sentence must be followed by the ids of the segments relevant to the sentence.
Table 12: Format explanations. Multi statement was used for GovReport, Single Statement for all other datasets.
QSP Scientific Article: {document} [End of Document]
Question: {question}
NQ Document: {document} [End of Document]
Question: {question}
EI Document: {document} [End of Document]
Question: {question}
WIC Document: {document} [End of Document]
Claim: {statement}
Additional Info: {additional_info}
CNLI Contract: {document} [End of Document]
Statement: {statement}
GR Document: {document} [End of Document]
{question}
Table 13: Formatting of instance specific input. See Table 10 for examples of task specific inputs.
QSP GR
AF1 ATT UF1 Avg RL ATT Avg
Inst before and after 42 75 14 43.77 22 43 32.60
Inst before 47 70 33 50.31 25 10 17.62
Inst after 40 60 20 40.16 21 14 17.85
Inst before, 1 example 53 67 43 54.21 26 36 30.97
Inst before, 2 examples 50 59 42 50.40 26 44 35.11
Inst before, 3 examples 50 66 67 60.86 28 54 41.08
Table 14: Prompt optimization results on GPT-3.5 under the citation approach. Inst before / after refers to the position of the instruction being before / after the task specific input. For complete instructions, see C Avg: Average of scores for one task.
Task Explanation You are given a Scientific Article and a Question. Answer the Question as concisely as you can, using a single phrase or sentence. If the Question cannot be answered based on the information in the Article, answer "unanswerable". If the question is a yes/no question, your answer should be "yes", "no", or "unanswerable". Do not provide any explanation. (If the question can be answered, provide one or several evidence paragraphs that can be used to verify the answer. Give as few paragraphs as possible.)
Format explanation Your reply must have the following format:
"<answer> [X] [Y]"
In your reply, replace <answer> with your solution to the task. Your reply must be followed by the ids of the relevant segments from the document.
Example 1 Scientific Article: Automated Hate Speech Detection and the Problem of Offensive Language
Abstract {omitted} [End of Document]
Question: What type of model do they train?
Example 2 {omitted}
Example 3 {omitted}
Instance specific input Scientific Article: Combining Thesaurus Knowledge and Probabilistic Topic Models
Abstract {omitted} [End of Document]
Question: Which domains do they explore?
Table 15: Example of the final prompt format used for citation on QASPER.
Input Claim
QASPER Question, response “The answer to the question ’Which domains do they explore?’ is ’news articles related to Islam and articles discussing Islam basics”’
Natural Questions Question, response “The answer to the question ’who won the 2017 ncaa mens basketball tournament?’ is ’North Carolina”’
"Evidence Inference Question, response “There was no significant difference between the effect of good motivation/capability and the effect of inadequate motivation/capability on quality of life.”
“Good motivation/capability [significantly increased/significantly decreased] quality of life compared to inadequate motivation/capability”’
"Wice Claim, response “Having over 3,000 animals of nearly 400 different species, the zoo has slowly increased its visitors and now ranks as the number one outdoor tourist attraction in the state.”
“The claim ’Having over 3,000 animals of nearly 400 different species, the zoo has slowly increased its visitors and now ranks as the number one outdoor tourist attraction in the state’ is partially supported.”
"ContractNLI Claim, response “Receiving Party shall not reverse engineer any objects which embody Disclosing Party’s Confidential Information.”
“Receiving Party may reverse engineer any objects which embody Disclosing Party’s Confidential Information.”’
GovReport Response statement “Improper payments in Medicaid increased from \$29.1 billion in fiscal year 2015 to \$36.7 billion in fiscal year 2017.”
Table 16: Examples for claims constructed for attributability evaluation.
TRUE premise: {evidence} hypothesis: {claim}
Attrscore Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
### Instruction:
Verify whether a given reference can support the claim. Options: ""Attributable, Extrapolatory or Contradictory.
### Input:
Claim: {claim}
Reference: {evidence}
### Response:
Minicheck predict: {evidence}</s>{claim}
Table 17: Prompts for attributability evaluation models based on the respective publications.
#Params Reference
Gemma-7b-it 7B Team et al. (2024)
GritLM-7B 7B Muennighoff et al. (2024)
Mistral-7B-Instruct-v0.2 7B Jiang et al. (2023)
Vicuna-7B-v1.5-16K 7B Chiang et al. (2023)
Flan-T5-XL/XXL* 3B/11B Longpre et al. (2023)
LongChat-7B-v1.5-32K 7B Li et al. (2023a)
Llama3-8B-Instruct 8B See  [^23] https://llama.meta.com/llama3/
Table 18: Open source models considered in selection experiments. *: Flan-T5-XL was used in fine-tuning, Flan-T5-XXL was used in prompting experiments.
AF1 ATT UF1 Avg
Gemma 0.41 0.28 0.25 0.27
GritLM 0.42 0.30 0.22 0.26
Mistral 0.46 0.34 0.26 0.30
Vicuna 0.42 0.31 0.24 0.28
Flan-T5-XL 0.45 0.61 0.22 0.42
LongChat 0.42 0.29 0.24 0.26
Llama3 0.44 0.37 0.25 0.31
Table 19: results of Fine-tuning Model Selection
QASPER GovReport Avg
Model Answer F1 Attr Unans F1 R-L Attr
Gemma-7b-it 22 2 49 18 0 17
GritLM-7B 26 6 48 20 0 18
Mistral-7B-Instruct-v0.2 31 32 48 24 0 25
Vicuna-7B-v1.5-16K 21 12 49 21 2 19
Flan-T5-XXL 26 19 66 12 1 22
LongChat-7B-v1.5-32K 21 18 48 22 13 24
Llama3-8B 17 17 47 22 0 19
Table 20: Results of model selection for prompted open source models.
QASPER Natural Questions
AF1 ATT UF1 Avg AF1 ATT UF1 Avg
BM25 32 60 42 44.49 44 33 59 45.17
SBERT 29 69 39 45.56 47 53 60 53.36
Contriever 37 75 50 54.20 47 51 61 52.89
Dragon 39 79 52 56.65 48 48 63 52.93
GTR 36 70 46 50.54 46 48 61 51.47
Evidence Inference Wice
CF1 EF1 Avg - CF1 EF1 Avg -
BM25 77 16 46.46 - 29 36 32.71 -
SBERT 78 23 50.51 - 33 43 42.50 -
Contriever 83 29 55.71 - 26 42 33.87 -
Dragon 78 33 55.23 - 31 41 36.29 -
GTR 78 33 55.58 - 33 43 38.20 -
Contract NLI Govreport
CF1 EF1 Avg - RL ATT Avg
BM25 43 34 38.35 - 27 33 29.96 -
SBERT 42 35 38.21 - 28 36 31.79 -
Contriever 38 37 36.80 - 24 30 26.92 -
Dragon 39 37 37.64 - 21 19 20.16 -
GTR 44 39 41.39 - 23 37 30.00 -
Table 21: Retriever selection for retrieve-then-read. Retrievers were combined with GPT-3.5 and were evaluated on 100 dev instances per dataset. The retriever that resulted in the best average score was used in all further experiments for the respective combination of retrieve-then-read and task.
QASPER Natural Questions
AF1 ATT UF1 Avg AF1 ATT UF1 Avg
BM25 45 67 63 58.33 42 36 59 45.69
SBERT 40 65 51 51.79 41 39 55 45.13
Contriever 48 72 65 61.64 40 42 55 45.83
Dragon 49 73 56 59.20 42 42 58 46.95
GTR 47 71 63 60.17 42 41 58 46.63
Evidence Inference Wice
CF1 EF1 Avg - CF1 EF1 Avg -
BM25 83 49 66.12 - 44 61 52.38 -
SBERT 82 50 65.75 - 40 65 52.13 -
Contriever 83 57 70.42 - 36 60 48.14 -
Dragon 87 54 70.51 - 37 64 50.49 -
GTR 86 62 73.66 - 37 62 49.75 -
Contract NLI Govreport
CF1 EF1 Avg - RL ATT Avg
BM25 42 49 45.25 - 26 51 38.29 -
SBERT 44 56 50.11 - 27 60 43.38 -
Contriever 45 58 51.42 - 23 45 33.93 -
Dragon 52 55 53.29 - 22 43 32.12 -
GTR 46 54 49.94 - 23 51 37.23 -
Table 22: Retriever selection for reduced approaches. Retrievers were combined with GPT-3.5-reduced-citation and were evaluated on 100 dev instances per dataset. The retriever the resulted in the best average score was used in all further experiments for the respective combination of reduced-citation / reduced-post-hoc and task.
QASPER Natural Questions
AF1 ATT UF1 Avg AF1 ATT UF1 Avg
BM25 52 65 73 63.25 41 40 57 45.92
SBERT 52 55 73 64.12 41 50 57 53.35
Contriever 52 63 73 68.42 41 54 57 55.28
Dragon 52 69 73 71.11 41 54 57 55.28
GTR 52 59 73 66.27 41 54 57 55.28
Evidence Inference Wice
CF1 EF1 Avg - CF1 EF1 Avg -
BM25 86 25 55.60 - 86 25 55.60 -
SBERT 86 20 52.94 - 86 20 52.94 -
Contriever 86 27 56.60 - 86 27 56.60 -
Dragon 86 28 57.27 - 86 28 57.27 -
GTR 86 24 55.27 - 86 24 55.27 -
Contract NLI Govreport
CF1 EF1 Avg - RL ATT Avg
BM25 46 36 41.14 - 28 73 50.55 -
SBERT 46 36 41.24 - 28 60 44.14 -
Contriever 46 39 42.69 - 28 65 46.63 -
Dragon 46 40 43.16 - 28 68 48.22 -
GTR 46 41 43.57 - 28 61 44.61 -
Table 23: Retriever selection for post-hoc approaches. Retrievers were combined with GPT-3.5-post-hoc and were evaluated on 100 dev instances per dataset. The retriever the resulted in the best average score was used in all further experiments for the respective combination of post-hoc and task.
Input Query
QASPER Question, response “Which domains do they explore? news articles related to Islam and articles discussing Islam basics”
Natural Questions Question, response “who won the 2017 ncaa mens basketball tournament? North Carolina”
Evidence Inference Question, response “There was no significant difference between the effect of good motivation/capability and the effect of inadequate motivation/capability on quality of life.”
“Good motivation/capability [significantly increased/significantly decreased] quality of life compared to inadequate motivation/capability”
Wice Claim “Having over 3,000 animals of nearly 400 different species, the zoo has slowly increased its visitors and now ranks as the number one outdoor tourist attraction in the state”
ContractNLI Claim, response “Receiving Party shall not reverse engineer any objects which embody Disclosing Party’s Confidential Information.”
“Receiving Party may reverse engineer any objects which embody Disclosing Party’s Confidential Information.”
GovReport Response statement “Improper payments in Medicaid increased from \$29.1 billion in fiscal year 2015 to \$36.7 billion in fiscal year 2017.”
Table 24: Examples for queries for post-hoc evidence retrieval.
Input Query
QASPER Question “Which domains do they explore?”
Natural Questions Question “who won the 2017 ncaa mens basketball tournament?”
Evidence Inference Question “With respect to Quality of life, characterize the reported difference between patients receiving good motivation/capability and those receiving inadequate motivation/capability. Choose between ’significantly decreased’, ’no significant difference’, and ’significantly increased’.”
Wice Claim “Having over 3,000 animals of nearly 400 different species, the zoo has slowly increased its visitors and now ranks as the number one outdoor tourist attraction in the state”
ContractNLI Claim “Receiving Party shall not reverse engineer any objects which embody Disclosing Party’s Confidential Information.”
GovReport Paragraph “Medicaid has been on our high-risk list since 2003, in part, because of concerns about the adequacy of fiscal oversight and the program’s improper payments—including payments made…”
Table 25: Examples for queries for retrieve-then-read and reduced retrieval.
Figure 7: Annotation instructions for attributability model evaluation.
unanswerable
n/a
i don’t know
idk
not known
answer not in context
unknown
no answer
it is unknown
the answer is unknown
unavailable
not clear
i cannot provide
i cannot directly provide
i cannot answer
i cannot display
not clear
not available
not readily available
Table 26: Simple keyphrases to detect abstaining. See §G for details on how these were used.
Verbs Adverbs
provide explicitly
mention specifically
state directly
specify clearly
define
report
name
offer
Table 27: Verbs and adverbs that were combined to phrases to detect abstaining. See §G for details on how these were used.